# OmniNWM: Omniscient Driving Navigation World Models

### v3

@@ Abstract
Figure 1: Versatile capabilities of OmniNWM. (a) State: Given a conditional image and ego trajectory, OmniNWM jointly generates comprehensive multi-modal outputs, including panoramic RGB, semantic, metric depth, and 3D occupancy videos. (b) Action: Conditioned on different input trajectories, OmniNWM facilitates precise panoramic camera control by converting them into normalized Plücker ray-maps as pixel-level guidance. (c) Reward: OmniNWM enables unbounded long-term navigation through a closed-loop pipeline: the planning trajectory guides the multi-modal generation, while dense rewards are derived from the generated 3D semantic occupancy. Note that all the input trajectories are encoded into normalized Plücker ray-maps before guiding the generation process, even when not explicitly visualized.
@@ 2 Related Work
World Models for Autonomous Driving. Recent progress in driving world models has advanced visual realism and controllability [100, 18, 116, 42, 16, 44, 79, 8, 98, 90, 19, 99, 66, 109, 58, 34, 36, 49, 64, 45, 76]. Early efforts like DriveDreamer [78] leverage multi-stage pipelines to incorporate traffic rules. MagicDrive [17] introduces cross-view attention for street-view synthesis, while Vista [18] improves generation via latent replacement strategies. UniScene [45] proposes a hierarchical occupancy-centric framework. Other notable works include ReSim [93] for simulation data augmentation, HUGSIM [116] for 3D Gaussian Splatting-based rendering, Drive-WM [79] for multi-view forecasting, and GEM [23] for egocentric prediction. However, existing models typically lack unified multi-modal generation (e.g., semantic map, 3D occupancy)  [78, 17, 18], and integrated reward modeling for closed-loop evaluation [108, 16, 45]. Our work addresses these limitations with long-horizon, multi-modal generation, precise control, and intrinsic rewards.
Camera-controlled Video Generation. Camera control has become a critical technique for guiding generative models in dynamic 3D environments [24, 48, 25, 81, 39, 4, 95, 29]. CameraCtrl [24] introduces a plug-and-play module to parameterize camera trajectories. MotionCtrl [81] decouples camera and object motion for independent control, while CamCo [89] introduces epipolar constraints to improve geometric consistency. VD3D [1] embeds Plücker coordinates into transformer architectures for view synthesis, and ReCamMaster [4] focuses on re-rendering existing videos from novel camera paths. However, these methods focus on short-range, monocular sequences, limiting their applicability to large-scale multi-view environments in autonomous driving [24, 25, 81, 89, 1, 4]. OmniNWM extends camera controllability to panoramic, long-horizon settings through a normalized Plücker ray-map representation.
@@ 3.1.1 Pixel-level Aligned Multi-model Generation
Joint generation of pixel-aligned multi-modal outputs has shown great potential to enhance the representational capacity of generative models and downstream performance [54, 45]. In OmniNWM, we initialize modality-specific noise latents for RGB, semantics, and depth, then concatenate them channel-wise for joint denoising using the PDiT. A shared decoder produces spatially aligned RGB, semantic, and depth video outputs. To ensure cross-modal consistency during training, semantic maps are colorized before VAE encoding and converted back to discrete labels via nearest-neighbor matching after decoding. The details of data curation on semantic and depth labels are provided in the supplementary material. This joint generation scheme ensures strong pixel-level alignment across modalities, which is essential for generating high-quality 3D occupancy, as detailed in the following subsection.
Figure 3: Occupancy generation. The generated panoramic RGB images, semantics, and depths are used to produce a 3D voxel volume. This design enables lightweight occupancy prediction compared with prior volumetric-based methods [45, 111], while naturally supporting the integration of occupancy-grounded rewards.
@@
where $\otimes$ denotes the outer product. $\texttt{Adap}_{d}$ and $\texttt{Adap}_{s}$ are adaptive aggregation networks for depth and semantic contexts. The resulting volume feature $\widehat{\mathbf{V}}$ is passed to a generation head with up-sampling and a softmax layer to produce the final 3D semantic occupancy. Unlike volumetric-based models like [45, 111] that directly synthesize full occupancy grids at high computational cost, our design computes occupancy from lightweight pixel-aligned features, which scales to long sequences and, crucially, enables intrinsic occupancy-grounded dense driving rewards.
@@
By transforming 3D occupancy into dense waypoint-level feedback, the proposed reward function enables fine-grained evaluation of policy behavior under diverse scenarios. In a critical oncoming-truck scenario (Figure 4), the reward function effectively distinguishes driving quality: a high-speed collision (left) yields the lowest reward with severe $R_{\text{col}}$ penalty; an insufficient avoidance (middle) reduces reward; and a successful evasion (right) achieves the highest reward with minimal control penalty $R_{\text{col}}$ and velocity penalty $R_{\text{vel}}$, confirming discriminative capability. More evaluation results can be found in the supplementary.
@@ 3.1.3 Trajectory Planning for Closed-loop Navigation
To fully leverage OmniNWM’s multi-modal outputs, we developed a semantic-geometric reasoning vision-language-action (VLA) model based on Qwen-2.5-VL [3], termed OmniNWM-VLA, as the default planning agent. We introduce a plug-and-play Tri-Modal Mamba-based Interpreter (Tri-MIDI) to integrate semantic, depth, and RGB contexts before tokenization and input into OmniNWM-VLA. To adapt planning trajectories suitable as Plücker ray-map inputs, we extend the output channels of OmniNWM-VLA to predict not only the x and y coordinates of waypoints, but also their heading angles. In contrast to prior methods [11, 103], which produce key frames at 2HZ, our method operates at 12HZ on the NuScenes dataset. Note that for the initial video clip, the trajectory is produced by first generating multi-modal results (i.e., RGB, depth, and semantics) using a static trajectory, which are then fed into OmniNWM-VLA to enable multi-modal reasoning for the planner. Moreover, OmniNWM supports straightforward integration of various planning agents for closed-loop evaluation. Additional experimental comparisons are provided in Section 4.2, while further architectural and implementation details of OmniNWM-VLA are available in the supplementary.
@@ 3.2.1 Efficient Plücker Ray-map Encoding
As illustrated in Figure 2, input trajectories are encoded into ray-maps via a parameter-free normalized Plücker encoder and injected into PDiT blocks. The normalized panoramic ray-maps are first downsampled in spatial and temporal dimensions to align with diffusion latents, and patchified into Plücker embedding tokens. These tokens are concatenated with the diffusion latent tokens (obtained via patchification of the diffusion latents) before being fed into the 3D full-attention layers within the PDiT blocks.
@@
${d}_{u,v}={R}{K}^{-1}[u,v,1]^{T}+{t}.$ (7)
The embedding of $i^{\text{th}}$ frame and $j^{\text{th}}$ view is represented as ${P}_{i,j}\in\mathbb{R}^{6\times h\times w}$, while a panoramic sequence of $n$ frames yields ${P}\in\mathbb{R}^{n\times 6\times v\times h\times w}$.
@@ 3.2.2 Scale and Pose Invariant Normalization
$\hat{{p}}^{(k)}_{u,v}=\left({o}_{k\to 0}\times\hat{{d}}^{(k\to 0)}_{u,v},\hat{{d}}^{(k\to 0)}_{u,v}\right)$. (11)
@@ 4 Experiments
This section presents evaluation results of our approach against SOTA approaches across multiple applications.
@@ 4.2 Main Results
Trajectory Planning Evaluation. Figure 7 (a) presents closed-loop evaluation results for trajectory planning, showing Scenario Pass Rate (SPR) [67] and average rewards calculated across 150 scenes from the NuScenes validation set. Impromptu-VLA [11] achieves higher performance compared to Qwen-2.5-VL [3], with results comparable to our OmniNWM-VLA. Figure 7 (b) illustrates the distribution of average rewards—evaluated using our occupancy-grounded driving rewards—for different VLA planning baselines. The horizontal axis represents the percentage of the number of evaluation scenes. The results confirm that our reward function effectively distinguishes between the planning trajectories generated by these VLA planners.
@@
Camera Control Accuracy. We evaluate camera control accuracy on the NuScenes validation set using Rotation Error (RotErr) and Translation Error (TransErr) metrics following previous works [24, 25]. Camera pose sequences are extracted using VGGT [75], with similarity scores calculated via Similarity Transformation (Sim(3)) [28]. As shown in Table 5 and Figure 8, our method with the panoramic normalized ray-map encoding significantly outperforms UniScene [45], yielding lower error distributions across the validation set. The RotErr achieved by our approach closely approaches ground-truth performance, while the higher TransErr reflects natural variability in driving distances across different scenarios. More evaluation results are provided in the supplementary.
@@ 5 Conclusion and Future Work
OmniNWM establishes a new paradigm for driving world models by unifying state, action, and reward in a single omniscient framework. It jointly generates long-horizon, pixel-aligned panoramic videos of RGB, semantics, depth, and 3D occupancy (state); enables precise, generalizable camera control via normalized Plücker ray-maps (action); and introduces occupancy-grounded dense rewards for closed-loop evaluation (reward). This integration delivers state-of-the-art performance in generation fidelity, control accuracy, and zero-shot generalization—making OmniNWM a powerful and holistic framework for autonomous driving development. Limitations and Future Work. While highly capable, the model remains computationally intensive, motivating future work on efficient deployment.
@@ A.1 Semantic Occupancy Prediction
Semantic Occupancy Prediction (SOP), also referred to as Semantic Scene Completion (SSC), is a holistic 3D perception task that jointly infers the semantic labels and geometric completeness of a scene from partial observations [86, 41, 43, 13, 46, 106, 7, 32, 107, 80]. Existing methods can be broadly categorized into LiDAR-based and camera-based approaches. LiDAR-based methods benefit from precise 3D measurements and have advanced significantly through architectures designed to handle sparse and irregular point clouds. For instance, PolarNet [106] leverages a polar BEV representation to address uneven point distribution, enabling efficient real-time inference. SalsaNext [13] improves segmentation via residual dilated convolutions and Bayesian uncertainty estimation. RangeNet++ [59] uses range images and a custom post-processing step to mitigate discretization artifacts. SCPNet [86] introduces multi-scale feature aggregation and dense-to-sparse distillation to enhance representation learning. Camera-based methods have gained traction due to their lower cost and broader applicability. MonoScene [7] pioneers monocular 3D semantic completion by projecting 2D features into 3D space. StereoScene [41] further bridges stereo geometry and BEV features for improved semantic and geometric consistency. Subsequent works such as TPVFormer [32] introduce tri-perspective views for richer 3D representations, while SurroundOcc [83] and OccFormer [107] leverage multi-view images and transformer-based aggregation for dense occupancy estimation. Despite these advances, most existing methods rely on limited modalities, hindering their robustness in complex driving scenarios. In contrast, our approach first generates pixel-aligned multi-modal outputs—RGB, semantic, and metric depth—from a reference image, then leverages these complementary modalities to predict 3D semantic occupancy. This joint generation strategy fully exploits both geometric and contextual information, leading to more accurate and consistent 3D scene understanding.
@@ A.2 Vision-Language Models
Recent years have witnessed growing interest in integrating Vision-Language Models (VLMs) and Large Language Models (LLMs) into autonomous driving systems, capitalizing on their strong reasoning capabilities and world knowledge [2, 110, 50, 65, 15, 91, 72, 113, 33, 51, 114, 103]. These models often exhibit emergent reasoning skills, such as chain-of-thought (CoT) reasoning, which can be harnessed for interpretable decision-making. For example, DriveVLM [72] combines LLM-based trajectory proposals with an end-to-end refinement module for final planning. SSR [51] addresses the limited spatial understanding of VLMs by converting depth maps into textual rationales and distilling them into lightweight embeddings for plug-and-play spatial reasoning. DriveGPT4 [91] uses LLMs in an iterative QA framework to explain driving behaviors and predict control signals. Doe-1 [113] reformulates driving as a next-token prediction task using a multi-modal generative model, while EMMA [33] encodes non-sensor inputs and outputs as natural language to fully leverage pre-trained LLMs. AutoVLA [118] utilizes raw visual inputs and language instructions for semantic reasoning and trajectory planning. FSDrive [103] introduces a pre-training paradigm for spatio-temporal reasoning, and Impromptu-VLA [11] provides a large-scale annotated dataset and baseline models for vision-language-action tasks in unstructured driving scenarios. In this work, we introduce OmniNWM-VLA, a semantic-geometric reasoning agent built upon Qwen-2.5-VL [2]. It integrates multi-modal inputs—RGB, depth, and semantic maps—via a Mamba-based [20] fusion block, enabling comprehensive scene understanding and trajectory planning.
@@ Appendix B Architecture Overview of OmniNWM-VLA
As shown in Figure 11, OmniNWM-VLA is a semantic-geometric reasoning agent designed to bridge high-level scene understanding with precise trajectory planning in autonomous driving. Built upon Qwen-2.5-VL [2], it integrates multi-modal inputs—RGB, depth, and semantic maps—through a novel Mamba-based [20] fusion block, enabling comprehensive spatial reasoning and actionable trajectory prediction. The agent outputs not only waypoint coordinates but also heading angles, which are essential for generating normalized Plücker ray-maps used in panoramic video generation. In this way, our model supports precise control through Plücker ray-maps, bridging the gap between high-level reasoning and low-level action execution in a unified framework.
@@
Tri-Modal Mamba-based Interpreter. At the core of OmniNWM-VLA is the Tri-Modal Mamba-based Interpreter (Tri-MIDI), a lightweight and plug-and-play module that fuses visual, geometric, and semantic information into a unified latent representation. This design allows the model to reason jointly over appearance (RGB), 3D structure (depth), and scene semantics (segmentation), which is critical for forecasting geometrically consistent and semantically plausible trajectories.
Given aligned multi-view inputs—RGB images $X_{V}$, metric depth maps $X_{D}$, and semantic segmentation maps $X_{S}$—we first stitch them into a unified panoramic grid $P_{V},P_{D},P_{S}$ to maintain spatial consistency across views. Each modality is encoded separately using pre-trained encoders: CLIP [62] for RGB, SigLIP [104] for depth, and SegFormer [88] for semantics:
@@
By integrating multi-modal perception with sequential reasoning, OmniNWM-VLA supports closed-loop evaluation and long-horizon planning, effectively connecting high-level decision-making with low-level action execution in a unified autonomous driving pipeline.

### v4

@@ Abstract
Figure 1: Versatile capabilities of OmniNWM. (a) State: Given a conditional image and ego trajectory, OmniNWM jointly generates comprehensive multi-modal outputs, including panoramic RGB, semantic, metric depth, and 3D occupancy videos. (b) Action: Conditioned on different input trajectories, OmniNWM facilitates precise panoramic camera control by converting them into normalized Plücker ray-maps as pixel-level guidance. (c) Reward: OmniNWM enables unbounded long-term navigation through a closed-loop pipeline: the planning trajectory guides the multi-modal generation, while dense rewards are derived from the generated 3D semantic occupancy. Note that all input trajectories are encoded into normalized Plücker ray-maps before guiding the generation process, even when not explicitly visualized.
@@ 2 Related Work
World Models for Autonomous Driving. Recent progress in driving world models has advanced visual realism and controllability [100, 18, 116, 42, 16, 44, 79, 8, 98, 90, 19, 99, 66, 109, 58, 34, 36, 49, 64, 45, 76]. Early efforts like DriveDreamer [78] leverage multi-stage pipelines to incorporate traffic rules. MagicDrive [17] introduces cross-view attention for street-view synthesis, while Vista [18] improves generation via latent replacement strategies. UniScene [45] proposes a hierarchical occupancy-centric framework. Other works include ReSim [93] for simulation data augmentation, HUGSIM [116] for 3D Gaussian Splatting-based rendering, Drive-WM [79] for multi-view forecasting, and GEM [23] for egocentric prediction. However, existing models typically lack unified multi-modal generation (e.g., semantic map, 3D occupancy)  [78, 17, 18], and integrated reward modeling for closed-loop evaluation [108, 16, 45]. Our work addresses these limitations with long-horizon, multi-modal generation, precise control, and intrinsic reward modeling.
Camera-controlled Video Generation. Camera control has become a critical technique for guiding generative models in dynamic 3D environments [24, 48, 25, 81, 39, 4, 95, 29, zhang2024cameras]. CameraCtrl [24] introduces a plug-and-play module to parameterize camera trajectories. MotionCtrl [81] decouples camera and object motion for independent control, while CamCo [89] introduces epipolar constraints to improve geometric consistency. VD3D [1] embeds Plücker coordinates into transformer architectures for view synthesis, and ReCamMaster [4] focuses on re-rendering existing videos from novel camera paths. However, these methods focus on short-range, monocular sequences, limiting their applicability to large-scale multi-view environments in autonomous driving [24, 25, 81, 89, 1, 4]. OmniNWM extends camera controllability to panoramic, long-horizon settings through a normalized Plücker ray-map representation.
@@ 3.1.1 Pixel-level Aligned Multi-model Generation
Joint generation of pixel-aligned multi-modal outputs has shown great potential to enhance the representational capacity of generative models and downstream task performance [54, 45]. In OmniNWM, we initialize modality-specific noise latents for RGB, semantics, and depth, then concatenate them channel-wise for joint denoising using the PDiT. A shared decoder produces spatially aligned RGB, semantic, and depth video outputs. To ensure cross-modal consistency during training, semantic maps are colorized before VAE encoding and converted back to discrete labels via nearest-neighbor matching after decoding. The details of data curation on semantic and depth labels are provided in the supplementary material. This joint generation scheme ensures strong pixel-level alignment across modalities, which is essential for generating high-quality 3D occupancy, as detailed in the following subsection.
Figure 3: Occupancy generation. The generated panoramic RGB images, semantics, and depth are used to produce a 3D voxel volume. This design enables lightweight occupancy prediction compared with prior volumetric-based methods [45, 111], while naturally supporting the integration of occupancy-grounded rewards.
@@
where $\otimes$ denotes the outer product. $\texttt{Adap}_{d}$ and $\texttt{Adap}_{s}$ are adaptive aggregation with SE3D for depth and semantic contexts. The resulting volume feature $\widehat{\mathbf{V}}$ is passed to a generation head with up-sampling and a softmax layer to produce the final 3D semantic occupancy. Unlike volumetric-based models [45, 111] that directly synthesize full occupancy grids at high computational cost, our design computes occupancy from lightweight pixel-aligned features, which scales to long sequences and, crucially, enables intrinsic occupancy-grounded dense driving rewards.
@@
By transforming 3D occupancy into dense waypoint-level feedback, the proposed reward function enables fine-grained evaluation of policy behavior under diverse scenarios. In a critical oncoming-truck scenario (Figure 4), the reward function effectively distinguishes driving quality: a high-speed collision (left) yields the lowest reward with severe $R_{\text{col}}$ penalty; an insufficient avoidance (middle) reduces reward; and a successful evasion (right) achieves the highest reward with minimal control penalty $R_{\text{col}}$ and velocity penalty $R_{\text{vel}}$, confirming discriminative capability. More evaluation results can be found in the supplementary material.
@@ 3.1.3 Trajectory Planning for Closed-loop Navigation
To fully leverage OmniNWM’s multi-modal outputs, we developed a semantic-geometric reasoning vision-language-action (VLA) model based on Qwen-2.5-VL [3], termed OmniNWM-VLA, as the default planning agent. We introduce a plug-and-play Tri-Modal Mamba-based Interpreter (Tri-MIDI) to integrate semantic, depth, and RGB contexts before tokenization and input into OmniNWM-VLA. To adapt planning trajectories suitable as Plücker ray-map inputs, we extend the output channels of OmniNWM-VLA to predict not only the x and y coordinates of waypoints, but also their heading angles. In contrast to prior methods [11, 103], which produce key frames at 2HZ, our method operates at 12HZ on the NuScenes dataset. Note that for the initial video clip, the trajectory is produced by first generating multi-modal results (i.e., RGB, depth, and semantics) using a static trajectory, which are then fed into OmniNWM-VLA to enable multi-modal reasoning for the planner.
Moreover, OmniNWM supports straightforward integration of various planning agents for closed-loop evaluation. Additional experimental comparisons are provided in Section 4.2, while further architectural and implementation details of OmniNWM-VLA are available in the supplementary.
@@ 3.2.1 Efficient Plücker Ray-map Encoding
As illustrated in Figure 2, input trajectories are encoded into ray-maps via a parameter-free normalized Plücker encoder and injected into PDiT blocks. Specifically, the normalized panoramic ray-maps are first downsampled in spatial and temporal dimensions to align with diffusion latents, and patchified into Plücker embedding tokens. These tokens are concatenated with the diffusion latent tokens (obtained via patchification of the diffusion latents) before being fed into the 3D full-attention layers within the PDiT blocks.
@@
${d}_{u,v}={R}{K}^{-1}[u,v,1]^{T}.$ (7)
The embedding of $i^{\text{th}}$ frame and $j^{\text{th}}$ view is represented as ${P}_{i,j}\in\mathbb{R}^{6\times h\times w}$, while a panoramic sequence of $n$ frames and $v$ views yields ${P}\in\mathbb{R}^{n\times 6\times v\times h\times w}$.
@@ 3.2.2 Scale and Pose Invariant Normalization
$\hat{{p}}^{(k)}_{u,v}=\left({o}^{(k\to 0)}\times\hat{{d}}^{(k\to 0)}_{u,v},\hat{{d}}^{(k\to 0)}_{u,v}\right)$. (11)
@@ 4 Experiments
This section presents evaluation results of our approach against existing approaches across multiple applications.
@@ 4.2 Main Results
Trajectory Planning Evaluation. Figure 7 (a) presents closed-loop evaluation results for trajectory planning, showing Scenario Pass Rate (SPR) [67] and average rewards calculated across 150 scenes from the NuScenes validation set. Impromptu-VLA [11] achieves higher performance compared to Qwen-2.5-VL [3], with results comparable to our OmniNWM-VLA. Figure 7 (b) illustrates the distribution of average rewards that are evaluated using our occupancy-grounded driving rewards for different VLA planning baselines. The horizontal axis represents the percentage of the number of evaluation scenes. The results confirm that our reward function effectively distinguishes between the planning trajectories generated by these VLA planners.
@@
Camera Control Accuracy. We evaluate camera control accuracy on the NuScenes validation set using Rotation Error (RotErr) and Translation Error (TransErr) metrics following previous works [24, 25]. Camera pose sequences are extracted using VGGT [75], with similarity scores calculated via Similarity Transformation (Sim(3)) [28]. As shown in Table 5 and Figure 8, our method with the panoramic normalized ray-map encoding significantly outperforms UniScene [45], yielding lower error distributions across the validation set. The RotErr achieved by our approach closely approaches ground-truth performance, while the higher TransErr reflects natural variability in driving distances across different scenarios. More evaluation results are provided in the supplementary material.
@@ 5 Conclusion and Future Work
OmniNWM establishes a new paradigm for driving world models by unifying state, action, and reward in a single omniscient framework. It jointly generates long-horizon, pixel-aligned panoramic videos of RGB, semantics, depth, and 3D occupancy (state); enables precise, generalizable camera control via normalized Plücker ray-maps (action); and introduces occupancy-grounded dense rewards for closed-loop evaluation (reward). This integration delivers state-of-the-art performance in generation fidelity, control accuracy, and zero-shot generalization, making OmniNWM a powerful and holistic framework for autonomous driving development. Limitations and Future Work. Despite its capabilities, the model remains computationally demanding, motivating future research into efficient deployment strategies.
@@ A.1 Semantic Occupancy Prediction
Semantic Occupancy Prediction (SOP), also referred to as Semantic Scene Completion (SSC), is a holistic 3D perception task that jointly infers the semantic labels and geometric completeness of a scene from partial observations [86, 41, 43, 13, 46, 106, 7, 32, 107, 80]. Existing methods can be broadly categorized into LiDAR-based and camera-based approaches. LiDAR-based methods benefit from precise 3D measurements and have advanced significantly through architectures designed to handle sparse and irregular point clouds. For instance, PolarNet [106] leverages a polar BEV representation to address uneven point distribution, enabling efficient real-time inference. SalsaNext [13] improves segmentation via residual dilated convolutions and Bayesian uncertainty estimation. RangeNet++ [59] uses range images and a custom post-processing step to mitigate discretization artifacts. SCPNet [86] introduces multi-scale feature aggregation and dense-to-sparse distillation to enhance representation learning. Camera-based methods have gained traction due to their lower cost and broader applicability. MonoScene [7] pioneers monocular 3D semantic completion by projecting 2D features into 3D space. StereoScene [41] further bridges stereo geometry and BEV features for improved semantic and geometric consistency. Subsequent works such as TPVFormer [32] introduce tri-perspective views for richer 3D representations, while SurroundOcc [83] and OccFormer [107] leverage multi-view images and transformer-based aggregation for dense occupancy estimation. Despite these advances, most existing methods rely on limited modalities, hindering their robustness in complex driving scenarios. In contrast, our approach first generates pixel-aligned multi-modal outputs of RGB, semantic, and metric depth from a reference image, then leverages these complementary modalities to predict 3D semantic occupancy. This joint generation strategy fully exploits both geometric and contextual information, leading to more accurate and consistent 3D scene understanding.
@@ A.2 Vision-Language Models
Recent years have witnessed growing interest in integrating Vision-Language Models (VLMs) and Large Language Models (LLMs) into autonomous driving systems, capitalizing on their strong reasoning capabilities and world knowledge [2, 110, 50, 65, 15, 91, 72, 113, 33, 51, 114, 103]. These models often exhibit emergent reasoning skills, such as chain-of-thought (CoT) reasoning, which can be harnessed for interpretable decision-making. For example, DriveVLM [72] combines LLM-based trajectory proposals with an end-to-end refinement module for final planning. SSR [51] addresses the limited spatial understanding of VLMs by converting depth maps into textual rationales and distilling them into lightweight embeddings for plug-and-play spatial reasoning. DriveGPT4 [91] uses LLMs in an iterative QA framework to explain driving behaviors and predict control signals. Doe-1 [113] reformulates driving as a next-token prediction task using a multi-modal generative model, while EMMA [33] encodes non-sensor inputs and outputs as natural language to fully leverage pre-trained LLMs. AutoVLA [118] utilizes raw visual inputs and language instructions for semantic reasoning and trajectory planning. FSDrive [103] introduces a pre-training paradigm for spatio-temporal reasoning, and Impromptu-VLA [11] provides a large-scale annotated dataset and baseline models for vision-language-action tasks in unstructured driving scenarios. In this work, we introduce OmniNWM-VLA, a semantic-geometric reasoning agent built upon Qwen-2.5-VL [2]. It integrates multi-modal inputs of RGB, depth, and semantic maps via a Mamba-based [20] fusion block, enabling comprehensive scene understanding and trajectory planning.
@@ Appendix B Architecture Overview of OmniNWM-VLA
As shown in Figure 11, OmniNWM-VLA is a semantic-geometric reasoning agent designed to bridge high-level scene understanding with precise trajectory planning in autonomous driving. Built upon Qwen-2.5-VL [2], it integrates multi-modal inputs of RGB, depth, and semantic maps through a Mamba-based [20] fusion block, enabling comprehensive spatial reasoning and trajectory prediction. The agent outputs not only waypoint coordinates but also heading angles, which are essential for generating normalized Plücker ray-maps used in panoramic video generation. In this way, our model supports precise control through Plücker ray-maps, bridging the gap between high-level reasoning and low-level action execution in a unified framework.
@@
Tri-Modal Mamba-based Interpreter. At the core of OmniNWM-VLA is the Tri-Modal Mamba-based Interpreter (Tri-MIDI), a plug-and-play module that fuses visual, geometric, and semantic information into a unified latent representation. This design allows the model to reason jointly over appearance (RGB), 3D structure (depth), and scene semantics (segmentation), which is critical for forecasting geometrically consistent and semantically plausible trajectories.
Given aligned multi-view inputs—RGB images $X_{V}$, metric depth maps $X_{D}$, and semantic segmentation maps $X_{S}$—we first embed them into latent tokens $P_{V},P_{D},P_{S}$ to preserve spatial consistency across views. Each modality is encoded separately using pre-trained encoders similar to the previous approach [51]. Specifically, we leverage CLIP [62] for RGB images, SigLIP [104] for depth maps, and SegFormer [88] for semantic maps:
@@
By integrating multi-modal perception with sequential reasoning, OmniNWM-VLA supports closed-loop evaluation and long-horizon planning, effectively connecting high-level decision-making with low-level action execution in a unified driving world model pipeline.